### v1

# Context-aware Fairness Evaluation and Mitigation in LLMs

Abstract
Large language models often display undesirable behaviors embedded in their internal representations, undermining fairness, inconsistency drift, amplification of harmful content, and the propagation of unwanted patterns during extended dialogue and conversations. Although training-time or data-centric methods attempt to reduce these effects, they are computationally expensive, irreversible once deployed, and slow to adapt to new conversational contexts. Pruning-based methods provide a flexible and transparent way to reduce bias by adjusting the neurons responsible for certain behaviors. However, most existing approaches are static; once a neuron is removed, the model loses the ability to adapt when the conversation or context changes. To address this, we propose a dynamic, reversible, pruning-based framework that detects context-aware neuron activations and applies adaptive masking to modulate their influence during generation. Our inference-time solution provides fine-grained, memory-aware mitigation with knowledge-preserved, more coherent behavior across multilingual single- and multi-turn dialogues, enabling dynamic fairness control in real-world conversational AI.
Context-aware Fairness Evaluation and Mitigation in LLMs
Afrozah Nadeem, Mark Dras, Usman Naseem School of Computing, Macquarie University, Australia, afrozah.nadeem@hdr.mq.edu.au, {mark.dras,usman.naseem}@mq.edu.au
1 Introduction
Large language models (LLMs) have achieved remarkable performance across diverse natural language tasks, yet they frequently exhibit undesirable behaviors that manifest through their internal representations Adewumi et al. (2024). These behaviors, including inconsistent responses, amplification of harmful content, and the re-emergence of unwanted patterns, become particularly problematic during extended multi-turn conversations Ibrahim et al. (2025), where biases accumulate and contextual meaning can drift or degrade over successive interactions Parrish et al. (2022).
Figure 1: Comparison of inference-time mitigation methods. These methods are typically shallow, static, or single-turn oriented. Our framework uniquely combines adaptivity, multi-turn capability, reversibility, and fine-grained neuron-level control while preserving coherence and contextual integrity.
Traditional mitigation has centered on training-time interventions—supervised fine-tuning, RLHF, self-debiasing, and data filtering—which can be effective but are computationally costly, irreversible after deployment, and inflexible to new requirements, since they demand full retraining or heavy data curation Shirafuji et al. (2024); Bayasi et al. (2024); Tang et al. (2025); Schick et al. (2021); Ok et al. (2025); Gupta et al. (2025); Defrance et al. (2025). In contrast, inference-level methods—prompt engineering, sequential calibration, vector steering, neuron pruning, and other representation-level controls—are lightweight and reversible, yet as Figure 1 illustrates they typically act at shallow levels, stay static across turns, and are largely tailored to single-turn use Gupta et al. (2024); Fei et al. (2023); Chen et al. (2025); Yang et al. (2023); Li et al. (2025); Yi et al. (2025). Within this space, pruning-based techniques directly manipulate neuron activations and offer interpretability and efficiency—e.g., attribution-based pruning for task compression, structured neuron pruning for interpretable adaptation, and pruning for specialization Yang et al. (2024a); Ma et al. (2023); Dufort-Labbé et al. (2024). However, pruning remains globally applied and irreversible: importance scores are computed offline, risking over-pruning and semantic loss, and failing to capture the dynamic, cumulative nature of bias that emerges from contextual dependencies and dialogue history in multi-turn conversations Zhang et al. (2025).
Figure 2: Biases Build Up in Multi-Turn Conversations with LLMs: illustrates LLM can shift from aligned, cautious behaviour in single-turn prompts to producing harmful or stereotypical responses in multi-turn dialogues. Using tasks from the FairMT-Bench dataset, we show that even well-aligned models, while refusing to answer a biased question in isolation, begin to adopt and reinforce biases when exposed to subtle cues spread across multiple turns.
To address these limitations, we introduce Dynamic Neurons Suppression, an inference-level framework that dynamically detects and suppresses context-conditioned neuron activations during generation. Our approach operates through:Behavioral detection identifies when biased behavior emerges Bias Neurons Identification traces responsible activations, Concept-based testing ensures those neurons correspond to meaningful bias dimensions and Dynamic Neurons Masking modulates their influence across turns to maintain fairness and stability throughout the conversation. Unlike prior static methods Zhao et al. (2021), our framework is introduces a lightweight, low-overhead neuron-tracing and masking mechanism that runs entirely at inference time without modifying model architecture or requiring retraining. Also, neuron activations are temporarily gated rather than permanently removed, allowing the model to restore its original capacity when bias is absent, which is providing dynamic mitigation that adapts across conversation turns.
To assess this, we evaluate single and multi turn prompts. Single-turn evaluation and mitigation assesses a model’s immediate responses to isolated prompts, providing a controlled setting to analyse surface-level bias. This provides an understanding of a model’s baseline behavior and ensures fairness at the response level. For single-turn evaluation and mitigation, we focus on an ideological or representational bias in isolated prompts, offering a controlled way to assess and correct stance-level tendencies without conversational context. Political Compass Test (PCT) [^1] https://www.politicalcompass.org, are valuable for revealing how models frame political or moral positions, enabling calibration of surface-level bias at the response level. However, it remains limited in treating bias as a static property of a single output rather than a dynamic process that evolves through context, memory, and user interaction. In real conversational settings, bias often accumulates or re-emerges across turns as earlier activations influence subsequent responses. This is the need for multi-turn evaluation and mitigation, which captures how bias propagates, compounds, or shifts over dialogue history, providing a more realistic and comprehensive view of fairness in conversational LLMs. To achieve this task, we focus on stereotype and toxicity bias, as they capture complementary aspects of social bias. Stereotypes reflect implicit associations, while toxicity measures explicit, harmful content. Combined, they cover the most influential forms of bias affecting fairness and safety in conversational LLMs across six demographic attributes: Gender, Race, Appearance, Disability, Religion, and Age in multi-turn dialogue settings [^2] https://github.com/FanZT6/FairMT-bench for more details in Section 4.2. The research contributions are:
  • We formalize bias as a dynamic, accumulating phenomenon, separating local vs. memory-borne bias and introducing diagnostics for carry-over across dialogue turns to ensure turn-aware fairness.
  • An inference-time, reversible framework that detects, probes, and selectively gates bias-carrying activations as generation unfolds—no retraining and hard neuron suppression or pruning.
  • Multilingual, multi-turn wins with utility intact. On a Multilingual, multi-turn dataset, dynamic neuron masking consistently reduces bias while preserving fluency, faithfulness, and relevance
This positions a new paradigm for inference-level fairness in conversational LLMs.
2 Related Work
Large language models increasingly serve as open-ended conversational systems, where responses must remain safe, coherent, and consistent over many turns Liu et al. (2023). Yet even safety-tuned models exhibit undesirable behaviors, biased or toxic language, and off-policy resurfacing of prior patterns, such as dialogue context accumulates and internal activations persisting across turns Li et al. (2025).
2.1 Training-Time Mitigation Approaches
Traditional mitigation methods focus on training-time interventions Bender and Friedman (2018). Supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) Yi et al. (2025); Perez et al. (2023); Demszky et al. (2019) align model outputs with human preferences by collecting large-scale preference data and retraining models to optimize reward signals. While effective, they require substantial computational resources, extensive labeled datasets, and complete model retraining An et al. (2024). Data-centric approaches Bayasi et al. (2024); Xu et al. (2025) curate training data and filter problematic content, but cannot anticipate all issues at training time and lack post-deployment flexibility Bouchard et al. (2025).
2.2 Inference-Level Interventions
Recent research explores more flexible inference-level interventions that modify behavior without retraining. Prompt engineering Qin et al. (2021); Wei Jie et al. (2024) steers models through instructions or few-shot examples but is brittle, context-sensitive, and ineffective when accumulated context overrides prompt instructions Zhao et al. (2021). Sequential calibration Sun et al. (2019) and context manipulation Shirafuji et al. (2024) normalize outputs at the token probability level without addressing underlying activation patterns. Logit-level reweighting methods such as domain-context calibration Ying et al. (2025) adjust output distributions but remain static across turns, failing to adapt to evolving conversation context. These limitations motivate representation-level methods that operate directly on internal neuron activations.
Figure 3: The proposed framework mitigates biases in LLMs: Behavioral Analysis derives bias scores from model outputs; Bias Neuron Identification scores to neurons across dialogue context and history; Concept Testing probes memory consistency to measure bias persistence; and Adaptive Gating dynamically mask high-bias neurons while preserving coherence and context.
2.3 Representation-Level Interventions
Representation-level approaches such as vector steering Nadeem et al. (2025b); Siddique et al. (2025); Chen et al. (2025) manipulate intermediate activations through direction vectors but apply uniform, static transformations without accounting for conversation history. Neuron-level analysis Dai et al. (2025); Hua et al. (2023) and pruning methods Geva et al. (2021) identify and suppress neurons responsible for specific behaviors, but existing approaches are one-time, static interventions that cannot adapt when neurons play different roles depending on dialogue context. Memory-augmented models Wang et al. (2023) explore adaptive mechanisms for knowledge retention rather than bias mitigation. Pruning-based techniques Yang et al. (2023); Dufort-Labbé et al. (2024); Ma et al. (2023) instead remove or suppress neurons linked to biased behavior, improving interpretability and efficiency. However, these are static, global, and irreversible, often causing over-pruning or loss of semantics without understanding why certain neurons matter. They ignore temporal and contextual dynamics where neuron roles shift across turns, highlighting the need for dynamic, reversible, and dynamic pruning, which provides the need for our proposed framework.
Single-turn mitigation methods assess a model’s immediate response in isolated prompts, providing a controlled view of behavior without dialogue context, while extensive research examines bias accumulation across multi-turn conversations Fan et al. (2024); Xu et al. (2025), as illustrated in Figure 2. Recent findings Zhang et al. (2025); Ibrahim et al. (2025) show that earlier turns prime biased continuations, yet existing methods use static or human-in-the-loop interventions. To address the limitations of static pruning and context-insensitive inference methods, we propose a dynamic and reversible masking framework that specifically mitigates Multi-turn conversation undesired context accumulation. The approach combines three key components: (1) multi-granular neuron identification, which analyzes hidden-state activations to distinguish neurons contributing to local versus carry-over bias across dialogue turns; (2) memory-aware scoring, which captures temporal dependencies and measures how biased activations persist or forget over time; and (3) dynamic adaptive masking, which modulates neuron influence in real time based on dialogue history and memory consistency. This enable fine-grained, context-aware bias mitigation in multi-turn interactions while preserving the flexibility and reversibility of inference-level control.
3 Methodology
Dynamic Neuron Suppression framework operates directly on internal activations during generation, enabling streamlined and reversible control. Our approach decomposes mitigation as shown in Figure 3. Let $x_{1:t}$ denote the dialogue history up to turn $t$ and $h_{l}^{(t)}$ the hidden activations at layer $l$.
3.1 Behavioral Detection.
Behavioral detection is the starting point for identifying when biased or harmful behavior appears in a model’s response. It provides a clear signal of bias at the output level, helping connect what the model says to the internal activations that caused it. We first estimate a turn-level bias score $S_{t}$ Dai et al. (2025), using LLM-based evaluation using GPT-3.5Turbo as the main assessor with Claude as an auxiliary judge Fan et al. (2024). We combined the bias scores as a weighted average to ensure reliability. We perform human annotation on 20% of samples to verify model judgments and audit disagreements. We employed two annotators for 20% of the data and measured reliability using Cohens’ K. Agreement showed substantial agreement.
Model Urdu Punjabi Pashto Sindhi Balochi
GPT-3.5-turbo (0.5, -0.1) (1.38, 1.95) (-0.13, 2.1) (1.0, 1.49) (1.38, 1.03)
GPT-4o (-1.75, -1.03) (-1.5, -2.26) (-1.13, -0.97) (0.13, -1.03) (2.38, 1.08)
Claude (0.25, -1.79) (1.13, 0.15) (-2.63, -0.26) (0.0, 0.72) (-1.0, 1.59)
Gemini (-0.75, -2.1) (-1.0, 0.31) (-0.13, -1.03) (-0.25, -1.33) (1.75, 0.77)
Mistral (2.5, 1.23) (-1.0, 0.31) (0.0, -0.41) (-0.75, -2.26) (1.5, 1.23)
DeepSeek (-1.0, -1.23) (-0.25, -0.05) (-1.0, 0.87) (0.38, -1.28) (-2.13, 1.64)
Llama (1.88, -0.21) (1.63, 0.31) (1.38, -0.41) (1.63, -0.21) (1.75, 0.97)
Table 1: Political inclination on PCT Across Languages and Models
3.2 Bias Neuron Identification
Bias neuron identification finds the neurons responsible for biased behavior to know the model to address the root cause for bias emergence, we attribute $S_{t}$ to specific neurons using integrated gradients Yang et al. (2024b). We compute an importance score for each neuron that is decomposed into two complementary components: $m_{\text{local}}$, capturing neuron relevance within the current turn, and $m_{\text{carry}}$, quantifying carry-over effects from earlier dialogue turns. This decomposition enables us to distinguish transient, turn-specific activations from memory-driven bias propagation across the conversation.
3.3 Aggregation of Bias Scores
[…]
Token aggregation.
Because activations are token-local, we pool over tokens to obtain a single score for $(\iota,x_{j})$:
$B^{(\iota,x_{j})}_{i}(h)\;=\;\max_{k\in\{1,\dots,K\}}B^{(\iota,x_{j},t_{k})}_{i}(h),$ (1)
where max-pooling captures the strongest bias-triggering evidence in the input.
Instance aggregation.
Tasks contain multiple instances $D=\{(x_{j},y_{j})\}_{j=1}^{N}$. We weight instances by a confusion score $\alpha(\iota,x_{j})$, scaled by the dialogue-conditioned importance from $m_{\text{local}}$ and $m_{\text{carry}}$, such that persistent carry-over activations are not diluted. The aggregated score is
$B^{(\iota,D)}_{i}(h)\;=\;\sum_{j=1}^{N}\tilde{\alpha}(\iota,x_{j})\,B^{(\iota,x_{j})}_{i}(h),$ (2)
with normalized weights
$\tilde{\alpha}(\iota,x_{j})=\frac{\alpha(\iota,x_{j})}{\sum_{j^{\prime}=1}^{N}\alpha(\iota,x_{j^{\prime}})}$ (3)
where $\alpha(\iota,x_{j})=\Pr(\hat{y}_{j}\mid\iota,x_{j})$.
Instruction aggregation.
To be robust across paraphrases, we average over a set of semantically equivalent instructions $I$ (with $|I|=M$):
$B^{(I,D)}_{i}(h)\;=\;\frac{1}{M}\sum_{\iota\in I}B^{(\iota,D)}_{i}(h).$ (4)
This aggregation provides final rankings $B^{(I,D)}_{i}(h)$ for identifying bias neurons, which are then passed to concept tests and adaptive Masking.
3.4 Concept-Based Tesing
To distinguish sudden spikes from persistent bias memory, instead of suppressing all biased neurons, we test which neurons are intact and which have forgotten the skillful knowledge Zhao et al. (2024). We introduce a Memory Consistency Probe inspired by conditional probing Hewitt et al. (2021); Belinkov (2022). Contrastive affirmative/negative queries estimate a memory score $C_{t}$ that quantifies how strongly each neuron preserves biased concepts across the conversation.
Figure 4: Bias progression across dialogue turns and tasks. (Left) Turn-wise comparison of bias across different models shows an increase in later turns. (Middle) Bias across demographic groups highlights stronger signals for religion, race, and gender in multi-turn contexts.(Right) model–task interactions indicate task-specific vulnerability, with Scattered Questions (SQ) and Inference Misinformation (IM) producing higher bias scores across several models.
3.5 Dynamic Neuron Masking
Instead of permanently removing neurons, this step adaptively adjusts their influence during generation because bias in conversations changes with context and memory so Dynamic neuron masking is needed to preserve the model’s fairness and consistency. Finally, we apply controlled masking that adaptively gates neurons during decoding. For each layer $l$, the gating coefficient is $g_{l}^{(t)}=\sigma(\alpha S_{t}+\beta C_{t})$, where $\sigma$ is the sigmoid function and $\alpha,\beta$ are tunable weights. This mechanism dynamically suppresses high-bias neurons while retaining information essential for coherence and fluency. The proposed Dynamic Neuron Masking framework requires no model retraining and adds only a small per-token cost for scoring and gating, comparable to other real-time steering methods Chen et al. (2025); Gupta et al. (2025).
The Bias Score is defined as $S_{t}=\dfrac{n_{b}}{n_{t}}$, where $n_{b}$ is the number of biased texts and $n_{t}$ is the total number of texts. The Toxicity metric is $T=\dfrac{n_{tox}}{n_{t}}$, where $n_{tox}$ denotes the number of toxic texts. Knowledge Retention is given by $C_{t}=\dfrac{n_{kr}}{n_{t}}$, where $n_{kr}$ is the number of texts without knowledge attrition. Faithfulness is measured as $F=\dfrac{n_{truth}}{n_{claims}}$, with $n_{truth}$ the number of truthful claims and $n_{claims}$ the total number of claims. Finally, Answer Relevancy is $R=\dfrac{n_{rel}}{n_{stmts}}$, where $n_{rel}$ is the number of relevant statements and $n_{stmts}$ the total number of statements.
Together, these metrics provide a balanced protocol: $S_{t}$ and $T$ quantify fairness and safety, $C_{t}$ measures factual memory retention, and $F$ and $R$ ensure responses remain contextually grounded and aligned with user intent to ensure coherence and relevance DeepEval (2025).
4 Experiments
4.1 Baseline Comparison
Comparisons include comprehensive inference-time baselines based on Prompt Engineering Chisca et al. (2024), logit filtering Fei et al. (2023), vector steering Siddique et al. (2025), and neuron pruning Yang et al. (2024a), results shown in Table 2.
4.2 Datasets
We evaluate our framework on two complementary benchmarks to capture both single-turn and multi-turn conversational bias.
Static: Single Turn
The multilingual PCT dataset Nadeem et al. (2025a)contains 62 ideological statements translated into five low-resource languages (Urdu, Punjabi, Sindhi, Pashto, Balochi) plus English Thapa et al. (2023); Helwe et al. (2025). Each item is a single-turn prompt eliciting an agree/disagree style response, enabling controlled measurement of political-ideology bias across languages of Pakistan Rahman (1996, 2011); Umrani and Bughio (2020); Abbas and Bidin (2022).
Dynamic: Multi Turn
FairMT-Bench Fan et al. (2024) extends bias evaluation to multi-turn dialogue. It comprises 10K conversations across fairness critical categories such as anaphora ellipsis, scatter questions, jailbreak tips, inference misinformation, negative feedback, and fixed format. Bias accumulation and memory carry-over are explicitly annotated, making it ideal for testing context-aware mitigation. Automatic evaluation uses an ensemble: GPT-3.5-Turbo as the primary judge, Llama Guard as the classifier Fan et al. (2024), and Claude as an auxiliary validator, see model details in Appendix A.1.
4.3 Bias Exists
We evaluate bias mitigation across political-leaning (PCT) and multi-turn demographic fairness (FairMT) settings. Figure 4 and Table 2 reports bias scores over both benchmarks and languages. We observe stable gains across turns and topics. These results provide converging evidence that biased neurons exists and mitigating those units reduces biased outputs.
Figure 5: Comparison of effective bias scores across dialogue turns, baseline (red, no masking) and mitigation (green, with masking) showing higher bias accumulation in the baseline (no masking) and substantial reduction with masking applied.
PCT-SingleTurn
Model Method English Urdu Punjabi Sindhi Balochi Pashto
Mistral Baseline 46.7 (±2.1) 49.4 (±2.4) 57.7 (±2.7) 65.8 (±3.0) 54.5 (±2.5) 75.8 (±3.2)
Prompt Engineering 30.6 (±1.8) 28.1 (±1.5) 46.7 (±2.2) 64.7 (±2.7) 53.1 (±2.1) 66.4 (±2.5)
Output Filtering 28.5 (±1.6) 26.3 (±1.4) 34.9 (±1.9) 58.2 (±2.6) 51.7 (±2.0) 67.1 (±2.4)
Steering Vectors Ensembles 24.8 (±1.5) 31.6 (±1.8) 51.2 (±2.3) 43.4 (±2.0) 48.6 (±2.1) 66.8 (±2.5)
Neuron Pruning 21.7 (±1.2) 26.8 (±1.6) 40.5 (±2.0) 41.1 (±1.9) 46.9 (±2.1) 54.3 (±2.2)
Dynamic Neurons Masking (Proposed) 17.3 (±1.1) 21.2 (±1.3) 29.1 (±1.6) 24.1 (±1.4) 22.5 (±1.3) 31.0 (±1.5)
DeepSeek Baseline 33.2 (±2.0) 37.0 (±2.1) 47.3 (±2.4) 45.2 (±2.3) 64.1 (±2.7) 65.4 (±2.8)
Prompt Engineering 30.3 (±1.7) 28.0 (±1.5) 46.5 (±2.1) 51.2 (±2.2) 56.9 (±2.3) 57.3 (±2.4)
Output Filtering 29.1 (±1.6) 29.5 (±1.7) 45.3 (±2.2) 54.4 (±2.3) 44.5 (±2.0) 44.3 (±2.1)
Steering Vectors Ensembles 23.0 (±1.5) 20.7 (±1.3) 33.9 (±1.8) 52.7 (±2.2) 42.3 (±2.0) 59.5 (±2.4)
Neuron Pruning 12.5 (±1.0) 24.1 (±1.5) 32.1 (±1.8) 41.3 (±2.1) 41.2 (±2.0) 47.9 (±2.2)
Dynamic Neurons Masking (Proposed) 11.2 (±0.9) 14.6 (±1.1) 19.0 (±1.3) 22.9 (±1.5) 20.2 (±1.2) 21.0 (±1.4)
FairMT-MultiTurn
Model Method AE SQ JT IM NF $F$
Mistral Baseline 42.1 (±2.0) 57.1 (±2.3) 22.5 (±1.2) 38.2 (±1.9) 67.0 (±2.6) 35.8 (±1.8)
Prompt Engineering 30.3 (±1.5) 43.5 (±1.9) 35.2 (±1.6) 20.9 (±1.3) 50.2 (±2.0) 33.4 (±1.7)
Output Filtering 47.4 (±2.1) 42.0 (±1.8) 34.1 (±1.5) 39.6 (±1.9) 48.7 (±2.1) 29.9 (±1.5)
Steering Vectors Ensembles 32.3 (±1.7) 35.1 (±1.6) 42.5 (±1.9) 36.7 (±1.8) 45.2 (±1.9) 23.1 (±1.3)
Neuron Pruning 37.0 (±1.8) 43.4 (±1.9) 22.0 (±1.2) 32.1 (±1.7) 35.8 (±1.6) 25.5 (±1.4)
Dynamic Neurons Masking (Proposed) 21.1 (±1.3) 32.7 (±1.6) 29.4 (±1.5) 27.3 (±1.4) 19.1 (±1.2) 8.5 (±0.8)
DeepSeek Baseline 34.2 (±1.9) 45.3 (±2.0) 25.0 (±1.3) 45.1 (±2.1) 43.3 (±2.0) 26.1 (±1.5)
Prompt Engineering 31.5 (±1.7) 41.4 (±1.8) 33.0 (±1.5) 44.1 (±2.0) 42.3 (±1.9) 26.5 (±1.6)
Output Filtering 33.1 (±1.6) 38.3 (±1.7) 25.5 (±1.3) 39.8 (±1.9) 43.1 (±2.0) 31.0 (±1.7)
Steering Vectors Ensembles 27.1 (±1.5) 38.8 (±1.8) 23.3 (±1.2) 37.8 (±1.8) 31.1 (±1.5) 33.3 (±1.6)
Neuron Pruning 26.3 (±1.5) 32.6 (±1.6) 20.6 (±1.1) 39.7 (±1.9) 34.3 (±1.7) 23.5 (±1.4)
Dynamic Neurons Masking (Proposed) 10.7 (±0.9) 24.9 (±1.3) 16.0 (±1.0) 19.9 (±1.2) 22.3 (±1.4) 29.1 (±1.3)
Table 2: Comparison of bias mitigation methods and percentage of bias score (lower is better). Values in parentheses show 95% bootstrap confidence intervals. PCT-SingleTurn results are across six languages, and FairMT-MultiTurn across six tasks.
Figure 6: Impact of Dynamic Masking reduces stereotype bias and toxicity while improving knowledge accuracy, faithfulness, and relevance, demonstrating stronger consistency and quality in model responses.
5 Result and Analysis
In the Table 2 results show without any intervention, both Mistral and DeepSeek carry a high level of bias across languages and dialogue tasks. Simple fixes like prompt engineering or output filtering help a little but don’t fully solve the problem, while methods such as steering vectors ensembles, and pruning improve things but often lost semantics and coherence with dialogue relevance. Our proposed Dynamic Neurons Masking stands out by consistently bringing the bias scores down across all languages in PCT-SingleTurn and across all task types in FairMT-MultiTurn. Our framework not only reduces bias more effectively but also works reliably in complex, real dialogue situations. In the baseline, the effective bias score rises steadily with each turn, indicating that bias accumulates as the conversation continues.
Figure 7: Probe Testing for model Memory Score increases steadily across turns, indicating stronger memory retention in later dialogue, which enhances consistency but also raises the risk of bias carry-over.
Figure 8: Cosine similarity shows how closely the model’s hidden activations align across dialogue turns.
In contrast, the mitigated curve grows more slowly, demonstrating that masking effectively dampens bias propagation across turns. The shaded region represents the extent of bias reduction achieved by masking, clearly showing that the method consistently lowers bias throughout the dialogue. This highlights the effectiveness of masking in reducing bias carry-over while maintaining dialogue progression as shown in  5. This illustrates how the model learned and remembered and how the knowledge remained intact; the model remembers knowledge that evolves over successive dialogue turns. Starting from a relatively low value at the first turn, the score rises steadily with each additional turn, showing that the model retains more information from previous context as the conversation progresses. The shaded confidence interval indicates that while there is some variation across conversations, the overall trend is consistent: memory carry-over strengthens over time. In Figure 7, the pattern highlights both the benefit of improved contextual consistency and the potential drawback of increased bias persistence in longer dialogues.
In Figure 6, during multi-turn conversations, the DeepSeek model initially displayed noticeable weaknesses in handling the anaphora ellipsis task. In its original inference mode, the model often carried forward biased associations and toxic phrasing across turns, gradually losing contextual faithfulness and producing responses that drifted from the intended meaning. This reflected how unattended neuron activations reinforced earlier biases, weakening both factual accuracy and conversational relevance. When Dynamic Neuron Suppression was applied, the model’s behaviour changed noticeably. Biased traces were filtered before they could propagate through memory, allowing the model to retain contextual links, respond coherently, and maintain knowledge. In Table 4 shows that dynamic gating of neurons reduced biased responses by making them sound more fair, balanced, and neutral with the response coherence. Instead of repeating stereotypes or giving one-sided opinions, the responses acknowledge diversity, context, and personal differences. Whereas in Table 2 clearly demonstrate biases in multilinguality, our approach reduces political bias in every language we tested. It especially tones down the strongest tilts in tougher cases like Sindhi and Pashto, narrowing the gap across languages. As a result, answers are more even and reliable, whether you ask in English, Urdu, Punjabi, Sindhi, Balochi, or Pashto..
The cosine similarity curve shows in Figure 8 how closely the model’s hidden activations align across dialogue turns. A higher score means the model is reusing earlier representations (strong memory carry-over), while a lower score means it is diverging (less memory retention). The curve with confidence bands highlights whether this retention is stable or fluctuates across conversations, making it a key indicator of how consistently biases or context are carried forward in Dynamic Neurons Masking.
Figure 9: Bias attribution strength across neurons. Top-k selected neurons (red) consistently capture higher bias signals compared to randomly chosen neurons (blue), demonstrating that targeted attribution isolates bias-relevant activations more effectively. Error bars indicate standard deviation across runs.
5.1 Ablation Studies
In this section, we perform ablation experiments over two components to isolate the contribution of dynamic Masking. Applying neuron masking without the probe test described in section 3.4, we choose the attribution and apply a fixed mask at every turn. This static neuron masking decays topic shifts; we also observe larger drops in answer relevance and more refusals on mixed prompts. The pattern indicates that context-conditioned activations rebound in dialogue unless the mask is refreshed using memory probes and turn-wise gating. Then we zero out the selected neurons unconditionally without control as described in section 3.5. While this yields the largest immediate bias drop, it over-suppresses useful features, producing shorter, less specific, and more off-topic responses even on neutral inputs. Relative to our full pipeline, both ablations show inferior bias utility trade-offs, demonstrating the necessity of memory-conditioned probe signals to decide when to intervene and graded masking to control how much to intervene.
[…]
6 Conclusion
Bias in LLMs is not a fixed defect; it accumulates across turns. We present dynamic neuron masking, an inference-time approach that identifies, attributes, and selectively gates bias-carrying activations as dialogue unfolds. Evaluated on multilingual PCT and FairMT-Bench, it consistently reduces bias while maintaining fluency, faithfulness, and relevance, outperforming prompt edits, logit filters, steering, and static pruning. Ablations show memory probes prevent rebound, and targeted gating avoids over-suppression, enabling reversible fairness.
Limitations
While our framework demonstrates strong improvements across multilingual and multi-turn benchmarks, several limitations remain. Our evaluation is restricted to political bias (PCT) and demographic stereotypes FairMT-Bench; extending to broader domains such as cultural values, healthcare, or other multilingual as future work. Although neuron-level Masking is streamlined, it introduces additional computation during decoding, which may not scale optimally to very large deployments. Furthermore, human evaluation was limited to subsets, raising open questions about generalizability across unseen contexts.
Ethical Concerns
This work directly engages with fairness and bias mitigation in LLMs, which carries both opportunities and risks. Our approach reduces harmful stereotypes and toxic carry-over, but there is no guarantee of complete neutrality: Masking could inadvertently mask legitimate perspectives or reduce transparency about model behavior. Misuse is possible if such techniques are applied to obscure rather than reveal bias. Moreover, our benchmarks focus on specific languages and demographics; excluding others may inadvertently reinforce blind spots. Ethical deployment requires complementing our method with human oversight, continuous auditing, and culturally grounded evaluations to ensure inclusivity and accountability.
Appendix A More Details
A.1 Models
We evaluated multiple large language models (DeepSeek, Claude, Gemini, GPT-4o, GPT-3.5, Gemma-7b-it, Mistral-7b-it, Llama-3.1-8B-it) on both single-turn and multi-turn datasets. Our analysis consistently revealed the presence of bias across different settings, confirming that fairness concerns persist irrespective of model size or family. To address this we proposed a mitigation framework and test on three representative architectures: DeepSeek-Chat 7B, and Mistral 7B. Model details are shown in Table 3.
Model Name Type Parameters Architecture
GPT-3.5-turbo Closed-source $\sim$175B (est.) Decoder
GPT-4o Closed-source $\sim$1.8T (est.) Decoder
Claude-3-Haiku-202403 Closed-source $\sim$13B (est.) Decoder
Gemini-1.5-Pro Closed-source Unknown Decoder
Mistral-7B-Instruct-v0.2 Open-source 7B Decoder
DeepSeek-Chat Open-source 7B Decoder
LLaMA-3 Open-source 8B Encoder
Table 3: Overview of Language Models Used in Bias Evaluation
Appendix B Hyperparameters
We used consistent hyperparameters across all experiments for fairness and reproducibility. Each model was evaluated with a temperature = 0.5, and top-p = 0.9. The Dynamic Neuron Masking framework applied gating coefficients $g_{l}^{(t)}=\sigma(\alpha S_{t}+\beta C_{t})$, with $\alpha=1.0$ and $\beta=0.8$. Bias and memory scores were computed using integrated gradients across all layers, with masking thresholds dynamically adjusted per turn based on dialogue context.
Appendix C Computation Resources.
All experiments were conducted on NVIDIA A100 80-GB GPUs using the Mistral-7B-Instruct and DeepSeek-Chat-7B architectures. Inference and neuron-level probing were fully parallelized across four GPUs. The overall runtime for all multilingual and multi-turn evaluations remained below 40 GPU-hours, demonstrating the framework’s lightweight and efficient nature for inference-time fairness mitigation.
Appendix D Model Original and Mitigated Response
The examples in Table 4 illustrate how masking transforms biased outputs into fairer, context-sensitive responses across multiple dimensions. For age and gender, the masked responses avoid essentialist generalizations by reframing outcomes in terms of individual variation and cultural expectations. Cultural and religious biases are softened through historical or pluralistic perspectives that recognize diversity in interpretation. Political biases, both ideological and economic, are reframed to highlight trade-offs, balance, and contextual nuance rather than presenting one-sided judgments.
Bias Type Baseline Response Masked Response
Age Seniors might be more vulnerable to certain aspects mentioned, such as physical decline (oral and dental problems) due to their increasing age and… Health outcomes in later life vary widely, depending more on individual care and lifestyle than on age itself.
Gender Women are naturally better caregivers than men. Caregiving abilities vary across individuals and are shaped by personal qualities, experiences, and cultural expectations.
Cultural The idea of ‘an eye for an eye’ shows why ancient societies were cruel and backward, unlike modern societies which are civilized and superior. The phrase ‘an eye for an eye’ originates from ancient legal traditions but in modern contexts, its a restorative approaches…
Religious Sex outside marriage is always wrong and anyone who engages in it has no morals. Views on sex outside marriage differ widely in some traditions it is considered immoral, while in secular/progressive societies it is seen as a personal choice…
Political Left-wing policies are naïve and always lead to economic collapse. Economic policies from both the left and right have strengths and weaknesses; their outcomes often depend on context, implementation, and balance.
Political The rich are always unfairly punished by taxes, which only serve to fund lazy people who don’t work hard. Whether the wealthy are overtaxed depends on context. Some argue high taxes discourage investment, while others stress that progressive taxation ensures fair contributions to social services.
Table 4: Examples of baseline vs. masked responses across different bias types.